An electronics hobbyist is building a radio set to receive the AM band, with frequencies from520 kHz to 1500 kHz . The antenna, which also serves as the inductor in an LC circuit, has an inductance of 274 μH . She needs to add a variable capacitor whose capacitance she can adjust to tune the radio.

What is the minimum capacitance the capacitor must have?

Cmin = ?
Answer #1

use the formulaa t resonance W^2 = 1/LC

w = 2 pi f

where L is Indiuctance

C is capacitance

so C = 1/W^2L = 1/4pi^2 f^2 L

as C and f are inversly proportional

Cmin be at f max


so when f = 1500 K Hz

C = 1/(4*3.14 * 3.14 * 1500000*1500000*275e-6)

C min = 41 pF